# Stability theory over toroidal or Novikov-type base and canonical modifications

### v2

@@ 3.2. Higher $\Theta$-stable reduction theorem
Let $\mathbbm{k}$ be a field, $(R,\mathfrak{m},v)$ be an equicharacteristic divisorial valuation ring which is $\mathbbm{k}$-algebra with $R/\mathfrak{m}\simeq\mathbbm{k}$ (divisorial valuation $\mathbbm{k}$-algebra), or more generally a valuation $\mathbbm{k}$-algebra which satisfies Lemma 3.5, and $\mathcal{M}$ be an algebraic stack over $\mathbbm{k}$ with affine diagonal.
As in Notation 1, we set $T=\mathbb{G}_{m}\otimes_{\mathbb{Z}}N$, where $N$ is a free abelian group of rank $r$, and fix a rational regular polyhedral cone $\tau\subset N_{\mathbb{R}}$ and $(0\neq)\xi\in\tau$ which is irrational if $r>1$.
Consider an arbitrary morphism $f\colon\mathop{\mathrm{Spec}}\nolimits(R)\to\mathcal{M}$. Suppose that $f(\mathfrak{m})$ sits in a higher $\Theta$-stratum $\mathcal{Z}^{+}$ in $\mathcal{M}$ while $f(\eta_{R})$ is not in $\mathcal{Z}^{+}$, where $\eta_{R}$ denotes the generic point of $R$. Now we fix an isomorphism $\Theta^{r}\simeq\Theta_{\tau}$ for some rational polyhedral cone $\tau$ of a lattice $N$ of rank $r$, and fix some $\xi\in\tau\setminus N\otimes\mathbb{Q}$. Then, after
@@
we have the following canonical extension:
  1. (1)
    $f$ extends to a unique (for fixed $\tau$) extends to its elementary ($\tau$-)modification $\overline{f}\colon{\overline{\rm ST}}_{R}(\tau^{\prime})\to\mathcal{M}$ which satisfy the generic point $\eta_{V(\tau^{\prime})}$ of $V(\tau^{\prime})$ maps to outside $Z^{+}$ and there is a $T$-equivariant morphism from the N-spectrum $\iota\colon\Delta_{-\xi}\to\mathcal{M}$ which satisfies that
    • $\overline{f}(\eta_{V(\tau^{\prime})})\notin Z^{+},$
@@
  2. (2)
    (Canonicity of degenerations) Such $\iota$ is unique, up to translation by $T(\mathbbm{k})\simeq(\mathbbm{k}^{*})^{r}$.
  3. (3)
@@
Corollary 3.10.
In the setup of Theorem 3.8, we obtain a canonical modification $f|_{\xi^{\prime}}\colon\overline{\Delta_{\xi}}(R^{\prime})\to\mathcal{M}$ for $\xi^{\prime}=(\xi,c)\in\tau^{\prime}$ (cf., Definition 2.3) of $f\colon\mathop{\mathrm{Spec}}\nolimits(R)\to\mathcal{M}$ such that
@@
  • the closed point $c_{\xi}$ maps to outside $\mathcal{Z}^{+}$,
  • $\lim_{t\to\infty}\lambda_{\xi^{\prime\prime}}(t)\cdot f|_{\xi^{\prime}}(c_{\xi})\in\mathcal{Z}^{+}$ for any $\xi^{\prime\prime}\in N\otimes\mathbb{Q}$ which approximates $\xi$ enough (as (3)).
In particular, we obtain a canonical modification $f_{{\rm new},\xi}\colon\mathop{\mathrm{Spec}}\nolimits R_{\rm new}(\xi)\to\mathcal{M}$ which maps the closed point to outside $\mathcal{Z}^{+}$ (recall Lemma 3.5).
@@
As [HL14, Lemma 4.4.6] (cf., also [AHHLR22]) shows, there is an affine $\mathbbm{k}$-scheme $M$, algebraic $\mathbbm{k}$-split torus $T^{\prime}$ and a smooth surjective representable morphism $[M/T^{\prime}]\to\mathcal{M}$ which induces ${\rm Filt}^{r}([M/T^{\prime}])\to{\rm Filt}^{r}(\mathcal{M})$, at the level of $\mathbb{Z}^{r}$-filtration stack (as algebraic stacks by [AHR20, 5.10, 5.11], [AHR19, 6.23]) for any $r$. Hence, we reduce the proof to the case when $\mathcal{M}$ is a quotient stack of the form $M$ an algebraic scheme over $\mathbbm{k}$ acted by an algebraic $\mathbbm{k}$-split torus $T^{\prime}$ as above. We assume it henceforth.
From our assumption on the structure of higher $\Theta$-strata on $Z^{+}$ and the Białynicki-Birula type theorem [HL14, 1.4.7], there is a $T$-equivariant morphism $Z^{+}\times U_{\tau}\to M$, following Notation 1, for some polyhedral cone $\tau$. Since there is also the (compatible) $T$-action on $M$, they induce a $T$-equivariant rational map $U_{\tau}\times\mathop{\mathrm{Spec}}\nolimits(R)\dashrightarrow M$. From the condition that $Z^{+}$ gives a union of connected components in the mapping stack, it follows that $U_{\tau}\times\mathop{\mathrm{Spec}}\nolimits(R)\dashrightarrow M$ is indeed not defined in the neighborhood of $(p_{\tau},\mathfrak{m})$, hence we want to resolve the indeterminancy as follows.
Lemma 3.11.
@@
  • the corresponding toroidal blow up of $U_{\tau}\times\mathop{\mathrm{Spec}}\nolimits(R)$, which we denote as $\mathcal{U}_{\Sigma^{\prime}}$, there is a $T$-equivariant morphism $\tilde{f}\colon\mathcal{U}_{\Sigma^{\prime}}\to M$ which extends $\tilde{f}$.
  • There is some $c^{\prime}\in\mathbb{R}_{>0}$ such that $\mathbb{R}_{\geq 0}\xi^{\prime}\subset N_{\mathbb{R}}\oplus\mathbb{R}$ with $\xi^{\prime}=(\xi,c^{\prime})$ lies inside a $r$-dimensional rational polyhedral cone $\tau^{\prime}$ of $\Sigma^{\prime}$ and the convex hull $\hat{\tau}^{\prime}:=\tau^{\prime}+\mathbb{R}_{\geq 0}(\vec{0},1)$ is a cone inside $\Sigma^{\prime}$. Further, $\tilde{f}(V({\tau^{\prime}}),0)\notin Z^{+}$, where $V(\tau^{\prime})$ denotes the strata of $\mathcal{U}_{\Sigma^{\prime}}$ which corresponds to $\tau^{\prime}$ as in usual toric geometry.
@@
We consider $T$-equivariant local coordinates $z_{1},\cdots,z_{N}$ around the image of the $T$-invariant point $p_{\tau}\in U_{\sigma}$ and consider its pullback (as rational functions) to $U_{\tau}\times\mathop{\mathrm{Spec}}\nolimits(R)$ and denote them as $\frac{f_{i}}{g_{i}}$ with $T$-semiinvariant holomorphic (algebraic) $f_{i}$s and $g_{i}$s. Consider the ideal $\mathcal{J}:=\langle\{{\rm lcm}\{g_{i}\},f_{1}\frac{{\rm lcm}\{g_{i}\}}{g_{1}},\cdots,f_{N}\frac{{\rm lcm}\{g_{i}\}}{g_{N}}\}\rangle$ which can be decomposed as
$\mathcal{J}=\sum I_{\vec{m}}z^{\vec{m}},$
@@
The blow up of $\mathcal{J}$ resolves the indeterminancy of the rational map $U_{\tau}\times\mathop{\mathrm{Spec}}\nolimits R\to M$ from the definition, and it is “toroidal” as there is some subdivision of the rational polyhedral cone $\tau\oplus\mathbb{R}_{\geq 0}\subset N_{\mathbb{R}}\oplus\mathbb{R}$ which realizes it. We can and do assume henceforth that $\Sigma^{\prime}$ is a regular simplicial fan, by taking subdivision each time if necessary.
@@
Suppose that the rays of $\Sigma^{\prime}|_{\mathbb{R}_{\geq 0}\xi+\mathbb{R}_{\geq 0}(\vec{0},1)}$ are spanned by $(\xi,0)$ and $\{(\xi,c_{i})\mid i=1,2,\cdots,l\}$ with $c_{1}>c_{2}>\cdots>c_{l}$ and denote the $r$-dimensional cone which includes $(\xi,c_{i})$ as $\tau^{\prime}_{i}$ and which includes $(\xi,0)$ as $\tau_{l+1}=\tau^{\prime}_{l+1}$. By further subdividing $\Sigma^{\prime}$ if necessary, we can and do assume that all $\tau^{\prime}_{i}$ projects down to the same cone $\tau$ in $N_{\mathbb{R}}$. By shrinking $\tau$ and subdividing $\Sigma^{\prime}$ further if necessary, we can and do assume that $\{\tau^{\prime}_{i}\}_{i}\cup\{\tau\times\{0\}\}$ are the only $r$-dimensional cones (walls) in $\Sigma^{\prime}$.
Now, we take a cone of $\Sigma^{\prime}$ which includes $(\xi,[c_{i}-\epsilon,c_{i}])$ for $0<\epsilon\ll 1$ as $\tilde{\tau_{i}}$ for each $i=1,\cdots,l$. If $\tilde{f}(V(\tilde{\tau_{l}}))\in Z^{+}$, then from the assumption that $\mathcal{Z}^{+}$ gives a union of connected components in ${\rm Map}(\Theta_{\tau},\mathcal{M})$, it follows that $\tilde{f}(V(\tau_{l+1})\times\mathop{\mathrm{Spec}}\nolimits(R))\in Z^{+}$ so that $f(\eta_{R})\in Z^{+}$ which gives a contradiction.
@@
Therefore, there is some $i\in\{1,\cdots,l\}$ such that
  • $\tilde{f}(V(\tau_{i}))\notin Z^{+}$,
  • For any cone $\sigma$ of $\Sigma^{\prime}$ which intersects with $(\xi,c_{i}+\mathbb{R}_{>0})$, the corresponding strata $V(\sigma)$ of $\mathcal{U}_{\Sigma^{\prime}}$ maps into $Z^{+}$ i.e., $\tilde{f}(V(\sigma))\subset Z^{+}.$
Suppose $i>1$, we do some modification as follows. We take any $j\in\{1,\cdots,i-1\}$. If $\tilde{f}(V(\tau^{\prime}_{j}))\neq\tilde{f}(V(\tau^{\prime}_{j}))$, then take the cone $\tau^{\prime}_{j,j+1}$ of $\Sigma^{\prime}$ which contains $\tau^{\prime}_{j}+\tau^{\prime}_{j+1}$ (the Minkowski sum) consider a point $z\in\tilde{f}(V(\tau^{\prime}_{j,j+1}))\subset Z^{+}$. Then the existence of two degenerating families from $z$ via $T$-action along $U_{\tau}$ and $U_{-\tau}$ respectively, to each $T$-invariant point, contradicts again to the higher $\Theta$-strata condition of $\mathcal{Z}^{+}$. Indeed, as both families degenerate to the isotrivial families of $\tilde{f}(V(\tau_{j-1}))$ so that it contradicts with the condition of higher $\Theta$-strata. Therefore, $\tilde{f}(V(\tau^{\prime}_{j}))=\tilde{f}(V(\tau^{\prime}_{j}))$ for any $j\in\{1,\cdots,i-1\}$. Therefore, if we do coarsening of $\Sigma^{\prime}$ to remove the walls $\tau^{\prime}_{1},\cdots,\tau^{\prime}_{i-1}$, the morphism $\mathcal{U}_{\Sigma^{\prime}}\to M$ still descend as a morphism. Therefore, we can and do assume $i=1$ henceforth and simply write $c_{1}$ as $c^{\prime}$ and $\tau^{\prime}_{1}$ as $\tau^{\prime}$. We complete the proof of Lemma 3.11. ∎
We take a simplicial subcone $\sum_{i=1}^{r}\mathbb{R}_{\geq 0}v^{\prime}_{i}$ of $\tau^{\prime}$, where $\{v^{\prime}_{i}\}_{i}\cup(\vec{0},1)$ is a basis of of $N$, which still contains $\xi^{\prime}$. We also take the cone which contains $(\xi,[c^{\prime}-\epsilon,c^{\prime}])$ for $0<\epsilon\ll 1$ and denote it as $\widetilde{\tau}^{\prime}$. Write $v^{\prime}_{i}=(v_{i},\frac{l_{i}}{m_{i}})$ with $l_{i},m_{i}\in\mathbb{Z}_{>0}$ and set $m:={\rm lcm}\{m_{i}\}$.
@@
Take an extension of $R$ as follows. If $R$ is a discrete valuation ring, then consider $R\to R^{\prime}$ with ramification degree $m$, which is obtained as a localization of $R[x]/(x^{m}-\pi)$ at $(x)$. Note that $\{\frac{m_{i}v^{\prime}_{i}}{m}\}_{i}$s together with $(\vec{0},\frac{1}{m})$ gives a basis of a regular cone of $N\oplus\frac{1}{m}\mathbb{Z}$. From the construction, (base change of) $\tilde{f}$ is extendable to $\widetilde{\rm ST}_{R^{\prime}}(\tau^{\prime})\to M$ and ${\overline{\rm ST}}_{R^{\prime}}(\tau^{\prime})\to\mathcal{M}$ and the desired properties are satisfied from the construction and finish the proof of (1). Note that the extension $R\to R^{\prime}$ does not change when we shrink $\tau^{\prime}$, as it follows from the construction. If $R$ is not divisorial valution ring, then we instead apply Lemma 3.5.

### v3

@@ 3.2. Higher $\Theta$-semistable reduction theorem
Let $\mathbbm{k}$ be a field, $(R,\mathfrak{m},v)$ be an equicharacteristic divisorial valuation ring which is $\mathbbm{k}$-algebra with $R/\mathfrak{m}\simeq\mathbbm{k}$ (divisorial valuation $\mathbbm{k}$-algebra), or more generally a valuation $\mathbbm{k}$-algebra which satisfies Lemma 3.5, and $\mathcal{M}$ be an algebraic stack over $\mathbbm{k}$ which is a quotient stack $[M/G]$ where $M$ is an algebraic scheme over $\mathbbm{k}$ and $G$ is an algebraic affine group acting on $M$.
On the other hand, as in Notation 1, we set $T=\mathbb{G}_{m}\otimes_{\mathbb{Z}}N$, where $N$ is a free abelian group of rank $r$, and fix a rational polyhedral cone $\tau\subset N_{\mathbb{R}}$ and $(0\neq)\xi\in\tau$ which is irrational if $r>1$.
Consider an arbitrary morphism $f\colon\mathop{\mathrm{Spec}}\nolimits(R)\to\mathcal{M}$. Suppose that $f(\mathfrak{m})$ sits in a higher $\Theta$-stratum $\mathcal{Z}^{+}$ in $\mathcal{M}$ while $f(\eta_{R})$ is not in $\mathcal{Z}^{+}$, where $\eta_{R}$ denotes the generic point of $R$. Now we fix some $\xi\in\tau\setminus N\otimes\mathbb{Q}$. Then, after
@@
there exists the following canonical extension $\overline{f}$ of $f$:
  1. (i)
    $f$ extends to its unique (for fixed $\tau$) extended elementary $\tau$-modification (see Definition 3.4)
    (7) $\overline{f}\colon{\overline{\rm ST}}_{R}(\tau^{\prime})\to\mathcal{M}$
    which satisfy the following (note that it is an extension of $f$):
    The generic point $\eta_{V(\tau^{\prime})}$ of the one dimensional locus $V(\tau^{\prime})$ (recall from Definition 3.1) maps to outside $Z^{+}$ and there is a $T$-equivariant morphism from the N-spectrum $\iota\colon\Delta_{-\xi}\to\mathcal{M}$ which satisfies that
    • $\overline{f}(\eta_{V(\tau^{\prime})})\notin Z^{+}$ (avoiding the higher $\Theta$-strata),
@@
  2. (ii)
@@
  3. (iii)
    (Canonicity of degenerations) Just in this item (iii), assume $\mathcal{M}$ is quotient stack by an algebraic split torus i.e., of the form $[M/T]$ where $M$ is an algebraic $\mathbbm{k}$-scheme and $T$ an algebraic split $\mathbbm{k}$-torus, and fix a lift of $f$ as $\tilde{f}\colon{\rm Spec}(R)\to M$. Then, the $\iota$ in (i), with admits some $\overline{f}$ which lifts to $\widetilde{\rm ST}_{R}(\tau^{\prime})$, is unique up to translation by $T(\mathbbm{k})\simeq(\mathbbm{k}^{*})^{r}$.
@@
By restricting the obtained extension $\bar{f}$ of $f$, we obtain the semistable reduction type statement, generalizing [AHLH23, §6].
Corollary 3.10 (Semistable reduction along higher $\Theta$-strata).
In the setup of Theorem 3.8, there exists a canonical modification $f|_{\xi^{\prime}}\colon\overline{\Delta_{\xi}}(R^{\prime})\to\mathcal{M}$ for $\xi^{\prime}=(\xi,c)\in\tau^{\prime}$ (cf., Definition 2.3) of $f\colon\mathop{\mathrm{Spec}}\nolimits(R)\to\mathcal{M}$ such that
@@
  • the closed point $c_{\xi}$ maps outside of $\mathcal{Z}^{+}$,
  • $\lim_{t\to\infty}\lambda_{\xi^{\prime\prime}}(t)\cdot f|_{\xi^{\prime}}(c_{\xi})\in\mathcal{Z}^{+}$ for any $\xi^{\prime\prime}\in N\otimes\mathbb{Q}$ which approximates $\xi$ enough (as (ii)).
In particular, we obtain a canonical modification
$f_{{\rm new},\xi}\colon\mathop{\mathrm{Spec}}\nolimits R_{\rm new}(\xi)\to\mathcal{M}$
which maps the closed point to outside $\mathcal{Z}^{+}$ (recall Lemma 3.5).
Here, the Novikov type valuation spectrum/ring $\overline{\Delta_{\xi}}(R^{\prime})$ and its complete version $R_{\rm new}(\xi)$ are used to play the role of the new base, as it is natural (or canonical) from the construction in the actual proof below. Indeed, if we are satisfied with extension $R^{\prime}$ of $R$ as discrete valuation ring for $r>1$ case, there are so many choices (as various smaller sub $\mathbbm{k}$-algebras of $R_{\rm new}(\xi)$ or their completions etc).
@@
Our proof of Theorem 3.8 is divided into steps. The core idea and arguments are in Step 1 with the key lemma 3.11, which may be one feature of this paper, and the rest is about how to reduce to it.
Step 1:
for quotient stack by an algbraic torus,
Step 2:
for general quotient stack (by an affine algberaic group)
Step 1.
First, we assume that $\mathcal{M}=[M/T]$ with a presentation $M$ and an algebraic split $\mathbbm{k}$-torus $T$ and prove the theorem in this case. Later, we essentially reduce to this case. From our assumption on the structure of higher $\Theta$-strata on $\mathcal{Z}^{+}$ and the Białynicki-Birula type theorem [HL14, 1.4.7], $\mathcal{Z}^{+}=[Z^{+}/T]$ for some closed subscheme $Z^{+}\subset M$ and there is a morphism $f\colon{\rm Spec}(R)\to M$ and $T$-equivariant morphism $Z^{+}\times U_{\tau}\to M$ which extends $f$, following Notation 1, for some rational polyhedral cone $\tau$. Since there is also the (compatible) $T$-action on $M$, they induce a $T$-equivariant rational map $U_{\tau}\times\mathop{\mathrm{Spec}}\nolimits(R)\dashrightarrow M$. From the condition that $Z^{+}$ gives a union of connected components in the mapping stack, it follows that $U_{\tau}\times\mathop{\mathrm{Spec}}\nolimits(R)\dashrightarrow M$ is indeed not defined in the neighborhood of $(p_{\tau},\mathfrak{m})$. We resolve its indeterminancy in an explicit “toroidal” manner as the follwing lemma.
This is our core for the proof of Theorem 3.8 and is of purely scheme-theoretic nature.
Lemma 3.11 (key lemma).
@@
[…]
  • There is some $c^{\prime}\in\mathbb{R}_{>0}$ such that $\mathbb{R}_{\geq 0}\xi^{\prime}\subset N_{\mathbb{R}}\oplus\mathbb{R}$ with $\xi^{\prime}=(\xi,c^{\prime})$ lies inside a $r$-dimensional rational polyhedral cone $\tau^{\prime}$ of $\Sigma^{\prime}$ and the convex hull
    $\hat{\tau}^{\prime}:=\tau^{\prime}+\mathbb{R}_{\geq 0}(\vec{0},1)$
    is a cone inside $\Sigma^{\prime}$. Further, $\tilde{f}(V({\tau^{\prime}}),0)\notin Z^{+}$, where $V(\tau^{\prime})$ denotes the strata of $\mathcal{U}_{\Sigma^{\prime}}$ which corresponds to $\tau^{\prime}$ as in usual toric geometry (hence, it is an image of a $\mathbb{G}_{m}$-torsor).
@@
We consider $T$-equivariant local coordinates $z_{1},\cdots,z_{N}$ around the image of the $T$-invariant point $p_{\tau}\in U_{\sigma}$ and consider its pullback (as rational functions) to $U_{\tau}\times\mathop{\mathrm{Spec}}\nolimits(R)$ and denote them as $\frac{f_{i}}{g_{i}}$ with $T$-semiinvariant holomorphic (algebraic) $f_{i}$s and $g_{i}$s. Consider the ideal $\mathcal{J}:=\langle\{{\rm lcm}\{g_{i}\},f_{1}\frac{{\rm lcm}\{g_{i}\}}{g_{1}},\cdots,f_{N}\frac{{\rm lcm}\{g_{i}\}}{g_{N}}\}\rangle$ which is $T$-invariant from the definition and its blow up resolves the indeterminancy of the rational map $\mathop{\mathrm{Spec}}\nolimits R\times U_{\tau}\to M$ from the definition. Here, ${\rm lcm}\{-\}$ stands for the least common multiple but any regular function on ${\rm Spec}(R)\times U_{\tau}$ which divides $g_{i}$s for all $i$ plays the role.
From the complete reducibility of $T$, $\mathcal{J}$ which can be decomposed as
(8) $\mathcal{J}=\sum I_{\vec{m}}z^{\vec{m}},$
@@
Note that a uniformizer $\pi$ of $R$ gives a formally étale morphism to an affine $\mathbbm{k}$-line 777this is where we use the equi-characteristic assumption of $R$, but we believe this is likely only for the simplicity (and geometric nature) of our description for now. A more appropriate algebraic translation of the arguments should directly extend to the mixed characteristics case. which we write $c_{\pi}\colon{\rm Spec}(R)\to\mathbb{A}^{1}_{\mathbbm{k}}$. From the form (8) of $\mathcal{J}$, there is a $T$-invariant ideal $J\subset\Gamma(\mathcal{O}_{U_{\tau}\times_{\mathbbm{k}}\mathbb{A}^{1}_{\mathbbm{k}}})$ whose pullback by $c_{\pi}$ coincides with $\mathcal{J}\subset\Gamma(\mathcal{O}_{U_{\tau}\times_{\mathbbm{k}}{\rm Spec}(R)})$. Because of the $T$-invariance of $J$, normalization of the blow up of $U_{\tau}\times\mathbb{A}^{1}$ along $J$ is a toric $\mathbbm{k}$-morphism with respect to the subdivision of $\tau\times\mathbb{R}_{\geq 0}$ to a rational polyhedral fan $\Sigma^{\prime}$ i.e.,
$b_{\Sigma^{\prime}}\colon T_{N}{\rm emb}(\Sigma^{\prime})\to\mathbb{A}^{1}_{\mathbbm{k}}\times_{\mathbbm{k}}U_{\tau},$
[…]
@@
We can and do assume henceforth that $\Sigma^{\prime}$ is a regular simplicial fan, by taking finer subdivision if necessary. Suppose that the rays of $\Sigma^{\prime}|_{\mathbb{R}_{\geq 0}\xi+\mathbb{R}_{\geq 0}(\vec{0},1)}$ are spanned by $(\xi,0)$ and $\{(\xi,c_{i})\mid i=1,2,\cdots,l\}$ with $c_{1}>c_{2}>\cdots>c_{l}$ and denote the $r$-dimensional rational polyhedral cone which includes $(\xi,c_{i})$ as $\tau^{\prime}_{i}$ for $i\leq l$, and which includes $(\xi,0)$ as $\tau_{l+1}=\tau^{\prime}_{l+1}$.
By further subdividing $\Sigma^{\prime}$ and taking its subfan if necessary, we can and do assume that all $\tau^{\prime}_{i}$ projects down to the same cone $\tau$ in $N_{\mathbb{R}}$ which is regular simplicial, its support stil contains $\xi$, and that $\{\tau^{\prime}_{i}\}_{i}\cup\{\tau\times\{0\}\}$ are the only $r$-dimensional cones (walls) in $\Sigma^{\prime}$, and we also set $\tau_{l+1}=\tau\times\{0\}$. Now, we take a $(r+1)$-dimensional rational polyhedral cone of $\Sigma^{\prime}$ which includes the sub-interval
$(\xi,[c_{i}-\epsilon,c_{i}])\subset N_{\mathbb{R}}\times\mathbb{R}_{\geq 0}$
for $0<\epsilon\ll 1$ as $\tilde{\tau_{i}}$ for each $i=1,\cdots,l$, and likewise, a $(r+1)$-dimensional rational polyhedral cone of $\Sigma^{\prime}$ which includes the sub-interval
$(\xi,[c_{1},c_{1}+\epsilon])\subset N_{\mathbb{R}}\times\mathbb{R}_{\geq 0}$
as $\tilde{\tau}^{\prime}_{0}$. Also, just as a notation in this paper, we denote each $T$-orbit of $T_{N}{\rm emb}(\Sigma^{\prime})$ which corresponds to a cone $\sigma\in\Sigma^{\prime}$ as $\overline{V}(\sigma)$ and its pullback to $\mathcal{U}_{\Sigma^{\prime}}$ as $V(\sigma)$.
Note that $V(\tilde{\tau_{l}})\in V(\tau_{l+1}=\tau\times\{0\})$ and the $T$-action on the $1$-dimensional locus $V(\tau_{l+1})$ is trivial and connected. Indeed, the toric blow up $b_{\Sigma^{\prime}}$ restricts to ${\overline{V}(\tau_{l+1})}(\simeq\mathbb{A}^{1}_{\mathbbm{k}})\to p_{\sigma}\times\mathbb{A}^{1}_{\mathbbm{k}}(\subset{\rm Spec}(R)\times U_{\sigma})$ as an $T\times\mathbb{G}_{m}$-equivariant isomorphism, where $p_{\sigma}$ denotes the unique $T$-invariant closed point of $U_{\sigma}$, but the $T$-action on both the source and the target are trivial. Hence its base change gives an isomorphism $V(\tau_{l+1})\overset{\simeq}{\to}{\rm Spec}(R)\times p_{\sigma}\simeq{\rm Spec}(R)$ and the $T$-action on its source and target are again trivial. Hence, we claim that for such one dimensional $V(\tau_{l+1})$, we easily confirmed:
Claim 1.
$V(\tau_{l+1})$ is connected and $V(\tilde{\tau}^{\prime}_{l})$ is in the closure of $V(\tau_{l+1})$. The $T$-actions on it and the image of $M$ are both trivial.
Now we suppose $\tilde{f}(V(\tilde{\tau_{l}}))\in Z^{+}$ and obtain the contradiction as follows. Intuitively, the arguments is by “contagion of $Z^{+}$” to the image of the generic point of ${\rm Spec}(R)$. Firstly, if $\tilde{f}(V(\tilde{\tau_{l}}))\in Z^{+}$, it would imply
(9) $\tilde{f}(V(\tau_{l+1}))\subset Z^{+},$
from the above Claim 1. Furthermore, note that there is a natural $T$-equivariant morphism (retraction) $U_{\tau}\times{\rm Spec}(R)\to V(\tau_{l+1})\simeq{\rm Spec}(R)$. Hence, from the assumption of Definition 3.3 of higher $\Theta$-strata, that $\mathcal{Z}^{+}$ gives a union of connected components in ${\rm Map}(\Theta_{\tau},\mathcal{M})$, the connectedness of $U_{\tau}$ implies that $\tilde{f}(\eta_{R})\in Z^{+}$, where $\eta_{R}$ is the generic point of ${\rm Spec}(R)$. Hence, we get contradiction and it follows that $\tilde{f}(V(\tilde{\tau_{l}}))\notin Z^{+}$. From the proof, it also follows that $\tilde{f}(V(\tau_{l+1}))\cap Z^{+}=\emptyset$. We summarize here.
Claim 2.
$\tilde{f}(V(\tilde{\tau_{l}}))\notin Z^{+}$ and $\tilde{f}(V(\tau_{l+1}))\cap Z^{+}=\emptyset$.
@@
On the other hand, note that $\tilde{f}(V(\tilde{\tau}^{\prime}_{0}))=f(\mathfrak{m})\in Z^{+}$ and that $\tilde{f}(V(\tau^{\prime}_{1}))\subset Z^{+}$. Therefore, there is some $i$ between them i.e., $i\in\{1,\cdots,l\}$ such that
  • $\tilde{f}(V(\tau^{\prime}_{i}))\subset Z^{+}$ and thus $\tilde{f}(V(\tilde{\tau}^{\prime}_{i}))\in Z^{+}$,
  • $\tilde{f}(V(\tilde{\tau}_{i-1}))\in Z^{+}$.
Suppose there is such $i$ as above with $i>1$. Note that the closure $\overline{V(\tau^{\prime}_{i})}\simeq\mathbb{P}^{1}_{\mathbbm{k}}$ is $V(\tilde{\tau}^{\prime}_{i})\cup V(\tau^{\prime}_{i})\cup V(\tilde{\tau}^{\prime}_{i-1})$. Hence, it follows that $\tilde{f}(\overline{V(\tau^{\prime}_{i})}\simeq\mathbb{P}^{1}_{\mathbbm{k}})\subset Z^{+}$.
Now we recall that $Z^{+}$ has the meaning of attracting locus towards $U_{\tau}$-direction, in the context of generalized versions of the Białynicki-Birula decomposition. The original statement in [BB73, 4.8] was for smooth $M$ (and $r=1$), but if $M$ is singular, it is later defined categorically ([HL14, §1.4, Proposition 1.4.1] for general $r$ and [Dri13, §0.2, §1.3, §1.4], [DG14, §1.4-§1.7]) as indeed it only maps to $M$ as an unramified monomorphism (cf. [HL14, Corollary 1.4.3, Remark 1.4.4], [DG14, 1.6.2, 1.7.4]) rather than an immersion. Nevertheless, from our assumption in Definition 3.3 that that the natural evaluation morphism ${\rm ev}_{(1,\cdots,1)}\colon\mathcal{Z}^{+}\to\mathcal{M}$ is a closed immersion, its base change $Z^{+}\to M$ is also a closed immersion in our case.  [^8] In the context of algebraic reconstruction of K-moduli of Fano varieties, this issue is sorted out in [BHLLX21, Proposition 7.2] by using (the arguments of) [BLZ22, 5.2]. In particular, the action morphism
$\alpha\colon\mathbb{G}_{m}\times Z^{+}$ $\to M,$
$(t,x)$ $\mapsto t\cdot x$
extends to $Z^{+}\times\mathbb{A}^{1}_{t}\to M$. If we restrict to $t=0$ case and further to
$\tilde{f}(\overline{V(\tau^{\prime}_{i})}\simeq\mathbb{P}^{1}_{\mathbbm{k}})\times\{0\}\to M,$
its open locus $V(\tau^{\prime}_{i})$ maps to a point $\tilde{f}(V(\tilde{\tau}^{\prime}_{i}))$, hence $\alpha$ contracts $\tilde{f}(\overline{V(\tau^{\prime}_{i})}\simeq\mathbb{P}^{1}_{\mathbbm{k}})$ to a point in $M$.
Note that removal of $\tau^{\prime}_{i}$ from $\Sigma^{\prime}$ to a coarser fan, which we denote as $\Sigma^{\prime}(i)$, corresponds to a blow down of $\overline{\bar{V}(\tau^{\prime}_{i})}\simeq\mathbb{P}^{1}_{\mathbbm{k}}$ in $T_{N}{\rm emb}(\Sigma^{\prime})$ and the above discusssion shows that $\mathcal{U}_{\Sigma^{\prime}}\to M$ descends to $\mathcal{U}_{\Sigma^{\prime}(i)}$. By repeating this, we obtain the following.
Claim 3.
There exists $\tilde{f}$ which satisfies
  1. ($\mathrm{i}$)
    $\tilde{f}(V(\tilde{\tau}^{\prime}_{i}))\not\subset Z^{+}$ for all $i=1,\cdots,l$,
  2. ($\mathrm{ii}$)
    but $\mathcal{U}_{\Sigma^{\prime}}\to M$ is still a morphism (not only a rational map).
Therefore, by setting $c_{1}$ as $c^{\prime}$ and $\tau^{\prime}_{1}$ as $\tau^{\prime}$, we complete the proof of Lemma 3.11. ∎
We continue the proof of Theorem 3.8. We write the simplicial subcone $\tau^{\prime}$ as $\sum_{i=1}^{r}\mathbb{R}_{\geq 0}v^{\prime}_{i}$, where $\{v^{\prime}_{i}\}_{i}\cup(\vec{0},1)$ are generators of the extremal rays of $\tau^{\prime}$. We denote $\tilde{\tau}_{1}$ (of the above proof of Lemma 3.11) as $\widetilde{\tau}^{\prime}$. Write $v^{\prime}_{i}=(v_{i},\frac{l_{i}}{m_{i}})$ with $l_{i},m_{i}\in\mathbb{Z}_{>0}$ and set $m:={\rm lcm}\{m_{i}\}$.
@@
Take an extension of $R$ as follows. If $R$ is a discrete valuation ring, then consider $R\to R^{\prime}$ with ramification degree $m$, which is obtained as a localization of $R[x]/(x^{m}-\pi)$ at $(x)$. Note that $\{\frac{m_{i}v^{\prime}_{i}}{m}\}_{i}$s together with $(\vec{0},\frac{1}{m})$ gives a basis of a regular cone of $N\oplus\frac{1}{m}\mathbb{Z}$. From the construction, (base change of) $\tilde{f}$ is extendable to $\widetilde{\rm ST}_{R^{\prime}}(\tau^{\prime})\to M$ and ${\overline{\rm ST}}_{R^{\prime}}(\tau^{\prime})\to\mathcal{M}$ and the desired properties are satisfied from the construction and finish the proof of (i). Note that the extension $R\to R^{\prime}$ does not change when we shrink $\tau^{\prime}$, as it follows from the construction. If $R$ is not divisorial valution ring, then we instead apply Lemma 3.5.